OpenStudy (anonymous):

True or False: (1/3)(3n + 3) = n + 1. Write the solution in lowest terms. 1/3 - (-3/4) = 4/12 1/3

OpenStudy (callisto):

(1/3)(3n + 3) = (1/3)(3n) + (1/3)(3) = ???

OpenStudy (anonymous):

true false

OpenStudy (callisto):

1/3 - (-3/4) = 1/3 + 3/4 = 1x4 / 3x4 + 3x3/4x3 = 4/12 + 9/12 = (4+9) /12 What would you get from here?

OpenStudy (anonymous):

1:Write the solution in lowest terms. 1/3 - (-3/4) = a: 4/12 b: 1 1/12 c: 1/3 2.True or False: (1/3)(3n + 3) = n + 1. this is the two diff problem...

OpenStudy (callisto):

I know.. First, (1/3)(3n + 3) = (1/3)(3n) + (1/3)(3) = what would you get from here?

OpenStudy (callisto):

That's the second problem

OpenStudy (anonymous):

true :D

OpenStudy (callisto):

Yes

OpenStudy (anonymous):

the first one i got 1.083 but idk the fraction for it

OpenStudy (callisto):

For the first problem 1/3 - (-3/4) = 1/3 + 3/4 = 1x4 / 3x4 + 3x3/4x3 = 4/12 + 9/12 = (4+9) /12 What would you get from here?

OpenStudy (anonymous):

b?

OpenStudy (callisto):

Yes :)

OpenStudy (anonymous):

woot wott

OpenStudy (callisto):

Nah? What's the problem?

OpenStudy (anonymous):

nt its correct got a 89% thoo . -.-